# An Exploratory Study on Using Large Language Models for Mutation Testing

### v1

@@ 3.3.1. Models
Model Type Model Training Data Time Release Time Size
Closed GPT-3.5-turbo 2021/09 2023/03 20B
GPT-4-turbo 2023/04 2023/06 110B
Open CodeLlama-13b-Instruct 2023/08 13B
StarChat-$\beta$-16b 2023/06 16B
@@ 3.3.2. Prompts
To design effective prompts for mutation generation, we follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experimentation. Figure 1 shows the default prompt template in our study.
@@
To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2021, 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Following the existing study on few-shot examples (Deng et al., 2024), we sample 6 of the bugs from QuixBugs.
To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all 6 examples. Table 3 shows these examples.
@@ 3.4. Evaluation Metrics
To assess the utility of the LLM-generated mutations, we should define a set of metrics that comprehensively evaluate them. Therefore, we designed several categories of metrics capturing different aspects as follows: Cost Metrics, Usability Metrics, and Behavior Metrics. Besides these metrics, we also additionally report the Mutation Count and Mutation Score. Note that both of them cannot directly reflect the quality of generated mutations. For example, a higher mutation score does not indicate a higher quality, which may also be caused by generating a large proportion of mutants that are easily killed.
@@
3.4.1. Cost Metrics
The Cost Metrics indicate time and economic costs in mutation generation, which contains the following two metrics.
@@ 3.4.1. Cost Metrics
Cost per 1K Mutations:  the average money cost in US dollars to generate 1k mutations. For GPT-3.5 and GPT-4, we calculate the average cost by API usage, while for the open-source models, we employ the costs of renting cloud servers.
@@ 3.4.3. Behavior Similarity with Real Bugs
Ochiai Coefficient:  approximates the semantic similarity by passing and failing test cases (Offutt and Untch, 2001), which is also widely used in the field of fault localization (Zou et al., 2019). Following the existing study (Ojdanic et al., 2023a), we employ the Ochiai Coefficient to calculate how a mutation resembles its corresponding real bug. More concretely, let $M$ and $B$ be a mutation and its corresponding real bug, and let $fT_{M}$ and $fT_{B}$ be the set of their failing test cases respectively. $M$’ coefficient is computed by the following formula:
@@ 3.5.1. Settings for Mutation Generation
3.5.2. Settings for RQ1-RQ2
In the first part of our study, we intend to qualitatively and comparatively evaluate the capability of LLMs in generating mutations. We adopt the default settings, which contain two models (i.e., GPT-3.5-Turbo and CodeLlama), and use the default prompt template, generating mutations on the universal set of the 405 bugs in Section 3.2.
To understand the performance of the LLMs, we adopt LEAM (Tian et al., 2022), $\mu$Bert (Degiovanni and Papadakis, 2022), PIT (Coles et al., 2016), and Major (Just et al., 2011) as baselines. Among these approaches, LEAM is the state-of-the-art mutation generation approach which employs a model trained from 13 million real Java bugs. While $\mu$Bert is based on Bert (Devlin et al., 2018), a masked pre-trained language model that can not conversational respond to human-like text and code.PIT (Coles et al., 2016) and Major (Just et al., 2011) are the popular traditional mutation testing tools with human-predefined simple mutation operators. Note that for PIT and Major, we employ all mutation operators.
3.5.3. Settings for RQ3-RQ4
In the second part of our study, we intend to explore how different prompts and different models impact the performance of mutation generation. Due to budget limitations, we sample a subset of 105 bugs from our dataset, which consists of 10 sampled bugs from each Defects4J project and all ConDefects bugs.
In RQ3, for the prompts exploration, we use the GPT-3.5-Turbo model and switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get 4 types of prompts, shown as follows:
@@ 3.5.3. Settings for RQ3-RQ4
In RQ4, for the LLMs exploration, we use the default prompt and investigate the performance of GPT-3.5-Turbo, GPT-4-Turbo, CodeLlama-13b-Instruct, and StarChat-$\beta$-16b.
@@ 3.5.4. Settings for RQ5
Generating non-compilable mutations incurs additional meaningless costs, so in the third part of our study, we study the root causes behind LLM generating non-compilable mutations. Specifically, we sample 384 non-compilable mutations of each mutation generation approach and analyze their reason rejected by the Java compiler and the features of their surrounding code context.
@@ 4. Evaluation Results
We performed our experimental analysis on two cloud servers, each of which is equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
Metric Category Metric GPT 3.5 CodeLlama-13b LEAM $\mu$BERT PIT Major
D4J CD D4J CD D4J CD D4J CD D4J CD D4J CD
Mut. Score 0.731 0.734 0.716 0.698 0.529 0.682
Mut. Count 351,557 303,166 671,737 141,031 1,032,392 314,694
Cost Avg. Gen. Time 1.79 9.06 3.06 2.23 0.017 0.083
\$ Cost per 1K Mut. 0.288 0.577 0.195
Usability Compilability Rate 61.7% 79.6% 75.8% 73.3% 33.8% 50.9% 21.9% 26.6% 98.3% 92.8%
Useless Mut. Ratio 11.6% 9.2% 41.0% 9.3% 0.5% 9.6% 1.8% 4.6% 0.0% 0.0%
Eq. Mut. Rate 2.3% 2.1% 1.3% 0.7% 1.6% 1.0% 2.1% 3.1% 0.5% 0.8%
Behavior Real Bug Detectability 0.967 0.867 0.906 0.667 0.767 0.489 0.785 0.378 0.511 0.489 0.916 0.689
Coupling Rate 0.416 0.625 0.398 0.612 0.267 0.318 0.409 0.581 0.133 0.593 0.342 0.479
Ochiai Coefficient 0.638 0.689 0.390 0.378 0.286 0.400 0.230 0.578 0.271 0.489 0.519 0.600
@@ 4.1.1. Cost
For each technique, we record the generation duration and the cost for API usage or rending GPU servers, shown as the Cost lines of Table 4. In addition, we also report the mutation count and mutation score, although they do not fully reflect the overall quality of the mutation set.
From the table, we can find that GPT-3.5 and CodeLlama-30b-Instruct generate 351,557 and 303,166 mutations111 Although GPT-3.5 and CodeLlama-13b are instructed to generate the same number of mutations, they do not consistently follow the instructions, thereby yielding different numbers of mutations., with a mutation score of 0.731 and 0.734, respectively. Traditional approaches PIT and Major yield the lowest mutation score among all the approaches.
In terms of cost, GPT-3.5 and CodeLlama-13b cost 1.79s and 9.06s to generate mutations, respectively. While LEAM, $\mu$Bert, PIT, and Major cost 3.06s, 2.23s, 0.017s, and 0.083s, respectively, indicating that previous approaches are faster than LLM-based approaches in generating mutations. In terms of economic cost, we only record the API usage cost of GPT-3.5 and the cloud server cost for running CodeLlama and LEAM, because the remaining approaches barely need GPU resources. The result indicates that renting cloud servers to run the open-source models costs more than using API for GPT-3.5.
Finding 1: Previous approaches cost less than LLM-based approaches in generating mutations. For example, in terms of mutation generation time, GPT-3.5 takes 1.79s to generate one mutation, CodeLlama-13b takes 9.06s, while PIT and Major cost 0.017s and 0.083s, respectively. Regarding economy cost, traditional approaches barely need GPU resources, but GPT-3.5 costs 0.288\$ to generate 1,000 mutations using API, and running CodeLlama-13b on a cloud server costs 0.577\$.
@@ 4.1.2. Mutation Usability
The usability of mutations (i.e., whether the mutations can be used to calculate mutation score) is affected by the capabilities of LLMs in correctly understanding the task and generating compilable, non-duplicate, and non-equivalent mutations. The results are shown as the Usability rows in Table 4. Due to the potential threat of data leakage, we represent the data of Defects4J and Condefects separately. Note that the mutations of PIT are in the form of Java bytecode and cannot be transformed back to the source level, so we skip PIT when answering this RQ.
Compilability Rate: We observe that GPT-3.5 achieves a compilability rate of 61.7% for Defects4J and 79.6% for ConDefects datasets. While CodeLlama-13b achieves a higher compilability rate of 75.8% for Defects4J and 73.3% for ConDefects. The rule-based approach Major achieves the highest compilabity rate (i.e., 98.3% on Defects4J and 92.8% on ConDefects). Note that though Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true)
Useless Mutation Rate: We also observe that LLMs generate a significant portion of duplicate and useless mutations. For example, for Defects4J, GPT-3.5 generates 11.6% useless mutations, accounting for 18.8% of all compilable mutations (i.e., 11.6%/61.7%). In contrast, LEAM, $\mu$Bert, and Major seldom generate duplicate mutations, which is likely due to these approaches being governed by strict grammar rules that limit the production of redundant mutations. For example, Major iteratively applies different mutation operators producing no redundant mutants.
@@
Dataset Type GPT-3.5 CL-13b LEAM $\mu$Bert Major
Defecs4J # All 351,332 302,941 671,342 140,753 314,527
# Sampled 384 384 384 384 384
# Eq. Mut. 9 5 6 8 2
ConDefects # All 225 225 395 278 167
# Sampled 143 143 196 162 123
# Eq. Mut. 3 1 2 5 1
Equivalent Mutation Rate: Equivalent mutations change the source code syntactically without affecting the program’s functionality, thereby affecting the accuracy of mutation score calculations. Due to the undecidable nature of identifying equivalent mutations, it is not feasible to design an algorithm that can completely detect them. Therefore, we sample a subset of mutations for each technique and manually assess whether they are equivalent. To maintain a 95% confidence level and a 5% margin of error, we sampled a fixed number of mutations from each approach’s output, to calculate the Equivalent Mutation Rate. Following the protocol of existing studies (Islam et al., 2019; Guo et al., 2024), two authors, each with over five years of Java programming experience, first received training on the equivalent mutants, and then independently labeled the mutations. The final Cohen’s Kappa coefficient score (Viera et al., 2005) is more than 90%, indicating a high level of agreement. Table 5 shows the number of sampled mutations and the number of identified equivalent mutations.
We again observe that Major performs the best in generating non-equivalent mutants. In particular, as shown in the bottom row of Usability section in Table 4, for GPT-3.5, 2.3% of mutations generated on Defects4J are equivalent, while the rate is 2.1% on ConDefects. CodeLlama-13b shows lower rates, with 1.3% and 0.7% on Defects4J and ConDefects, respectively. LEAM generates equivalent mutations at rates of 1.6% on Defects4J and 1.0% on ConDefects. $\mu$Bert produces 2.1% on Defects4J and 3.1% on ConDefects. Major achieves the lowest rates, with only 0.5% on Defects4J and 0.8% on ConDefects.
When comparing the two datasets Defects4J and ConDefects for the three usability aspects introduced above, we observe that the metrics vary between the Defects4J and ConDefects datasets for each mutation approach. This variation may be attributed to differences in the complexity and structure of bugs present in each dataset. Mutations in Defects4J may involve APIs and global variables that are not explicitly mentioned in prompts, making the context more complex. This complexity results in both metrics being inferior on Defects4J compared to ConDefects.
Finding 2: Major outperforms LLM-based and other learning-based methods in terms of compilability rate, useless mutation rate, and equivalent mutation rate. In particular, GPT-3.5 and CodeLlama-13b exhibit compilability rates of 61.67% and 75.86%, useless mutation rates of 11.60% and 40.96%, and 2.28% and 1.14% equivalent mutations, respectively. In contrast, Major excels with a 98.33% compilability rate, 0% useless mutations, and 0.59% equivalent mutants.
Figure 2. Number of types and examples of newly introduced AST Nodes from different mutation approaches.

### v2

@@ 3.3.1. Studied Models
Model Type Studied Model Base Model Training Data Time Release Time Size
Closed GPT-3.5-Turbo GPT 2021/09 2023/03
GPT-4o GPT 2023/10 2024/05
GPT-4o-Mini GPT 2023/10 2024/07
Open StarChat-$\beta$-16b StarCoder 2023/06 16B
CodeLlama-Instruct-13b Llama 2023/08 13B
DeepSeek-Coder-V2-236b DeepSeek 2023/09 2024/07 236B
@@ 3.3.2. Prompts
Figure 1 shows the default prompt template in our study. To design effective prompts for mutation generation, we follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments.
@@
To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our default prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 3.
@@ 3.4. Evaluation Metrics
To assess the utility of the LLM-generated mutations, we should define a set of metrics that comprehensively evaluate them. Therefore, we designed several categories of metrics capturing different aspects as follows: Cost Metric, Usability Metrics, and Behavior Metrics. Besides these metrics, we also additionally report the Mutation Count and Mutation Score. Note that both of them cannot directly reflect the quality of generated mutations. For example, a higher mutation score does not indicate a higher quality, which may also be caused by generating a large proportion of mutants that are easily killed.
@@
3.4.1. Cost Metric
The Cost Metric indicates time costs in mutation generation.
@@ 3.4.3. Behavior Similarity with Real Bugs
Ochiai Coefficient:  approximates the semantic similarity by passing and failing test cases (Offutt and Untch, 2001), which is also widely used in the field of fault localization. Following the existing study (Ojdanic et al., 2023a), we employ the Ochiai Coefficient to calculate how a mutation resembles its corresponding real bug. More concretely, let $M$ and $B$ be a mutation and its corresponding real bug, and let $fT_{M}$ and $fT_{B}$ be the set of their failing test cases respectively. $M$’ coefficient is computed by the following formula:
@@ 3.5.1. Settings for Mutation Generation
3.5.2. Settings for RQ1-RQ3
In the first part of our study, we intend to qualitatively and comparatively evaluate the capability of LLMs in generating mutations. We adopt the default settings, which contain all six LLMs, and use the default prompt template, generating mutations on the universal set of the 851 bugs in Section 3.2.
To understand the performance of the LLMs, we adopt LEAM (Tian et al., 2022), $\mu$Bert (Degiovanni and Papadakis, 2022), PIT (Coles et al., 2016), and Major (Just et al., 2011) as baselines. Among these approaches, LEAM is the state-of-the-art mutation generation approach which employs a model trained from 13 million real Java bugs. While $\mu$Bert is based on Bert (Devlin et al., 2018), a masked pre-trained language model that can not conversational respond to human-like text and code.PIT (Coles et al., 2016) and Major (Just et al., 2011) are the popular traditional mutation testing tools with human-predefined operators. Note that for PIT and Major, we employ all mutation operators.
3.5.3. Settings for RQ4
In the second part of our study, we intend to explore how different prompts and different models impact the performance of mutation generation. Due to budget limitations, we sample a subset of 306 bugs from our dataset, which consists of 10 sampled bugs from each Defects4J project and all ConDefects bugs.
In RQ4, for the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the open-source DeepSeek-Coder-V2-236b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get 4 types of prompts, shown as follows:
@@ 3.5.4. Settings for RQ5
Generating non-compilable mutations incurs additional meaningless economic and time costs, so in the third part of our study, we study the root causes behind LLM generating non-compilable mutations. Specifically, we sample 384 non-compilable mutations of each mutation generation approach and analyze their reason for being rejected by the Java compiler and the features of their surrounding code context.
@@ 4. Evaluation Results
We performed our experimental analysis on four rented cloud servers, each equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
@@ 4.1.1. Cost
For each technique, we record the generation duration, as shown in the Avg. Gen. Time row of Table 4. In addition, we report the mutation count and mutation score, although they do not fully reflect the overall quality of the mutation set.
From the table, we find that GPT-3.5, GPT-4o, GPT-4o-Mini, DeepSeek, StarChat, and CodeLlama generate 538338, 539715, 588803, 594711, 463248, and 457195 mutations, respectively222 Although the LLMs are instructed to generate the same number of mutations, they do not consistently follow the instructions, thereby yielding different numbers of mutations., with a mutation score ranging from 0.516 to 0.726. Overall, traditional approaches (e.g., LEAM and PIT) yield lower mutation scores among all the approaches.
In terms of time cost, the six LLMs cost 1.79s, 1.76s, 1.65s, 4.25s, 7.53s, and 9.06s to generate mutations, respectively. In comparison, LEAM, $\mu$Bert, PIT, and Major take 3.06s, 2.23s, 0.02s, and 0.08s, respectively. The rule-based approaches, PIT and Major, are the fastest, while LLMs show a competitive speed with small model-based approaches, LEAM and $\mu$Bert.
Finding 1 (RQ1): Rule-based approaches cost less time than LLMs in generating mutations. For example, GPT LLMs take fewer than 1.8s per mutation. In contrast, PIT and Major only require 0.02s and 0.08s, respectively, while small model-based approaches (i.e., LEAM and $\mu$Bert) take 2-3s.
Metric GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert PIT Major
Mut. Score 0.697 0.726 0.708 0.694 0.516 0.700 0.545 0.675 0.581 0.486
Mut. Count 538338 539715 588803 594711 463248 457195 890417 200896 2890433 346282
Avg. Gen. Time 1.79 1.76 1.65 4.25 7.53 9.06 3.06 2.33 0.02 0.08
Comp. Rate 60.2% 75.6% 73.6% 75.5% 11.1% 70.2% 35.0% 22.5% —— 98.6%
Useless Mut. Rate 10.9% 7.8% 6.7% 8.3% 8.9% 39.3% 1.0% 1.7% 0.0% 0.0%
Eq. Mut. Rate 2.2% 1.2% 1.7% 1.2% 1.8% 1.2% 1.3% 2.5% 0.0% 0.6%
Real Bug Detec. 91.7% 93.4% 93.4% 92.8% 47.1% 83.1% 71.7% 71.3% 51.3% 74.4%
Coupling Rate 41.4% 43.6% 44.1% 41.5% 29.4% 39.8% 28.7% 41.6% 14.0% 36.0%
Ochiai Coeff. 65.0% 68.5% 66.9% 61.8% 19.3% 40.8% 37.4% 31.6% 31.5% 44.4%
@@ 4.1.2. Mutation Usability
The usability of mutations (i.e., whether the mutations can be used to calculate mutation score) is affected by the capabilities of LLMs in correctly understanding the task and generating compilable, non-duplicate, and non-equivalent mutations. The results are shown as the Comp. Rate, Useless Mut. Rate, and Eq. Mut. Rate rows in Table 4. Since the mutation generation process of the PIT model is not visible and it only outputs mutations that pass the tests, we excluded it from the compilability rate calculation.
Compilability Rate: Four out of the six LLMs (i.e., GPT-4o, GPT-4o-Mini, DeepSeek, CodeLlama) achieve over 70%, with one exceeding 60% (i.e., GPT-3.5). In contrast, the small model-based approaches, LEAM and $\mu$Bert, have rates at 35.0% and 22.5%, respectively. Compared to small models, LLMs like GPTs and DeepSeek show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 40% higher than that of LEAM and $\mu$Bert. The rule-based approach Major significantly outperforms, with nearly 99% of its mutations successfully compiling. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true).
Useless Mutation Rate: We also observe that LLMs generate a significant portion of duplicate and useless mutations. For useless mutations, the percentage generated by LLMs ranges from a minimum of 6.7% (GPT-4o-Mini) to a maximum of 39.3% (CodeLlama), and five out of the six LLMs have a useless mutation rate below 11%. In contrast, all the traditional approaches rarely generate duplicate mutations, which is likely due to them being governed by strict grammar rules that limit the production of redundant mutations. For example, PIT and Major iteratively apply different mutation operators producing no redundant mutants.
@@
Dataset Type GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert PIT Major
D4J #All 537109 538509 587583 593481 462042 455634 888452 199347 2888641 345292
#Sampled 384 384 384 384 384 384 384 384 384 384
#Eq. Mut 9 3 5 4 8 5 6 8 0 2
CD #All 1229 1206 1220 1230 1206 1561 1965 1549 1792 990
#Sampled 293 292 293 293 292 309 322 308 317 277
#Eq. Mut 6 5 7 4 4 3 3 9 0 2
Equivalent Mutation Rate: Equivalent mutations change the source code syntactically without affecting the program’s functionality, thereby affecting the accuracy of mutation score calculations. Due to the undecidable nature of identifying equivalent mutations (Budd and Angluin, 1982), it is not feasible to design an algorithm that can completely detect them. Therefore, we sample a subset of mutations for each technique and manually assess whether they are equivalent. Following the common practice of sampling (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015), to maintain a 95% confidence level and a 5% margin of error, we sampled a fixed number of mutations from each approach’s output, to calculate the Equivalent Mutation Rate. Following the protocol of existing studies (Islam et al., 2019; Guo et al., 2024), two authors, each with over five years of Java programming experience, first received training on the equivalent mutants, and then independently labeled the mutations. The final Cohen’s Kappa coefficient score (Viera et al., 2005) is more than 90%, indicating a high level of agreement. All labeling results are available in our package for reproducibility (our, 2024), and Table 5 presents the sampling and labeling results.
As shown in the Eq. Mut. Rate row of Table 4, we again observe that the rule-based approaches, PIT and Major, perform the best in generating non-equivalent mutants, with an equivalent mutation rate of less than 1.0%. All six LLMs exhibit an equivalent mutation rate ranging from 1.2% to 2.2%, similar to LEAM and $\mu$Bert. Overall, the equivalent mutation rates for all approaches studied are relatively low and can be ignored during mutation testing. We further analyzed the mutations generated by PIT and Major, and found that they apply specific mutation operators to a statement (e.g., for a < b, only generating !(a < b), true, and false), significantly reducing the equivalent mutation rate.
Finding 2 (RQ1): Rule-based approaches, PIT and Major, outperform others in terms of compilability rate, useless mutation rate, and equivalent mutation rate. In particular, GPT-4o exhibits a compilability rate of 75.6%, a useless mutation rate of 7.8%, and an equivalent mutation rate of 1.2%. In contrast, Major excels with a compilability rate of 98.3%, a useless mutation rate of 0%, and an equivalent mutation rate of 0.6%.
Figure 2. Number of newly introduced AST nodes of the studied approaches and the few-shot examples.